Status Report

NASA ISS On-Orbit Status 3 July 2008

By SpaceRef Editor
July 4, 2008
Filed under , , ,
NASA ISS On-Orbit Status 3 July 2008

All ISS systems continue to function nominally, except those noted previously or below.

Crew Sleep Cycle: Wake/sleep cycle remains right-shifted (5:30am – 9:00pm EDT).

Continuing preparations for the two upcoming Orlan EVAs, CDR Volkov & FE-2 Chamitoff began their day with a 3-hr. training session in Soyuz TMA-12 to familiarize Gregory with spacecraft ops during his isolation in the Descent Module (SA) during the EVAs. [With Sergey supervising, Gregory’s hands-on training focused on equipment familiarization, working with Soyuz communications facilities, monitoring & interfacing with the Soyuz Neptun-ME console displays & controls, reviewing his tasks during his stay in the SA, assisting the CDR & FE-1 with leak checking, pressure equalization between Orbital Module (BO) and SA, BO/SA hatch opening, Orlan suit doffing if required, etc.]

Afterwards, Chamitoff & FE-1 Kononenko worked in the US Airlock (A/L) on EMU (Extravehicular Mobility Unit) #3004 to disconnect and remove its power harness for its subsequent use on an Orlan-M suit for EVA-20A. [The electrical harness, which normally powers the glove heaters and the EMU TV, is going to be used, along with a US REBA (Rechargeable EVA Battery Assembly), on Orlan #26 to provide real-time EVA video coverage via the US EMU TV and helmet lights. The harness will be installed on Orlan #26 next week during the EVA Dry-Run. EMU #3004 is not planned for use during Increment 17, 18, or ULF2 and will be brought home on 15A.]

Afterwards, Kononenko continued previously started work on Orlan #26, equipping it up with the electrical cable to provide power from the REBA to the US wireless TV ERCA (EMU RF Camera Assembly) in the BRTA-1M telemetry systems unit installed earlier (6/27).

Later, the crew conducted a 1.5-hr review of the airlock procedures for DC1 Docking Compartment egress/ingress, supported with ground specialist tagup via S-band. FE-2 Chamitoff also tagged up with EVA specialists at MCC-H to discuss his support activities.

As part of their standard fitness evaluation, Kononenko & Volkov undertook another session with the Russian MO-5 MedOps protocol of Cardiovascular Evaluation during Graded Exercises on the VELO cycle ergometer, assisting each other as CMO (Crew Medical Officer). (Last time done: 5/21). [The 50-min assessment, supported by ground specialist tagup via VHF and telemetry monitoring, uses the Gamma-1 ECG equipment with biomed harness, skin electrodes and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer’s instrumentation panels. For the graded exercise, the subject works the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each. Data output involves a kinetocardiogram, rheoplethysmogram, rheoencephalogram and a temporal pulsogram.]

Chamitoff did another sample rearrangement in the MELFI (Minus Eighty Degree Laboratory Freezer for the ISS), swapping the latest blood and urine samples in their box modules between two dewars of different temperature, started yesterday. [Today, sample vials (separated in 6 bags) were swapped between Dewar 2 and Dewar 3. Rationale: The uncrewed station configuration during the upcoming Russian spacewalks will require MELFI being powered off for an extended period. Gregory’s current configuration changes between dewars are intended to provide best possible protection of the science samples contained in them. Toward this aim, Greg is packing as much thermal mass (“coldness”) into Dewar 2 to maximize the cold volume hold time.]

In the FGB, after having set up the Matryoshka-R (RBO-3-2) radiation payload on 7/1 with a new memory card (#950), the FE-1 today checked the contents of the new card for temperature, quantity and total file size, prepared all data for downlink and reinstalled the ALC-950 PCMCIA (Portable Computer Memory Card International Adapter) card in the Spectrometer. [RBO-3-2 is using the ESA/RSC-Energia experiment ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS/ALC) with its Spectrometer (AST) and ALC equipment, temporarily located in the FGB on panel 429 (normally in DC1).]

With the TVIS treadmill roller bearings approaching their end-of-life, the FE-2 spent a few minutes on the weekly inspection of the rollers, checking the treadmill’s belt both left and right for any noticeable depressions due to seized or worn rollers.

The FE-1 completed the periodic (about twice a month) replenishing of the Elektron oxygen generator’s water supply for electrolysis, filling the KOV EDV container with water transferred on 6/30 from the ATV WDS (Water Delivery System) tank 1 to an empty EDV. [The 40-minute procedure is specially designed to prevent air bubbles larger than ~10 mm from getting into the BZh Liquid Unit where they could cause Elektron shutdown.]

Oleg also conducted today’s routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

Later, working from the Russian discretionary “time permitting” task list, Kononenko completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (FE-2), and VELO bike with bungee cord load trainer (CDR/MO-5, FE-1/MO-5). Before sleeptime tonight, Oleg is to transfer the exercise data file to the MEC laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~1:35pm EDT, the CDR and both FEs participated in a seven-minute live televised PAO interview event with KGPE-TV (Mike Scott) in Fresno, CA.

At ~3:25pm, the ISS crew is scheduled for their regular weekly tagup with the Lead Flight Director at JSC/MCC-H via S-band/audio. [Private S/G-2 (Space-to-Ground 2) phone patch via SSC (Station Support Computer)].

A new task added to Gregory Chamitoff’s voluntary “job jar” is the transfer of hardware (ESA payload spares) from ATV to COL (Columbus Orbital Laboratory) and to unpack/rearrange/stow the CTBs (Cargo Transfer Bags) in the Columbus MSSP (Mobile System Stowage) & FSSP (Fixed System Stowage) Provisions or in other CTBs.

WRM Update: An updated Water Recovery Management “cue card” was uplinked overnight for the crew’s reference. [The new card (17-0002P) lists 35 CWCs (Contingency Water Containers,~1365.2 L total) for the four types of water identified on board: technical water (650.6 L, for Elektron, flushing, hygiene, incl. 509.4 L non-usable water because of Wautersia bacteria), potable water (706.7 L, incl. 260.6 L currently on hold), condensate water (4.9 L), waste/EMU dump and other (7.9 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

FSL Troubleshooting Update: On 7/1, during his troubleshooting of the Fluid Science Laboratory in ESA’s Columbus module, Chamitoff replaced the halogen lamp that appeared to be malfunctioning in the FSL CEM (Central Experiment Module). Yesterday, FSL rack checkouts by ground teams at COL-CC/Oberpfaffenhofen showed that Greg’s work was successful and the FSL is fully functional. Further FSL commissioning is required prior to starting Geoflow payload operation, but Geoflow can be installed.

Russian TVM Issue: As reported by Moscow, Lane 3 of the Russian Terminal Computer (TVM) system is showing off-nominal performance. Under investigation.

EVA-20A Timeline Preview: Orlan EVA-20a by Volkov/EV1 & Kononenko/EV2 is scheduled to begin at on 7/10 ~2:18pm EDT (DC1 EV hatch open), to last an estimated 6 hrs. During the spacewalk, EV2 will be riding on the DC1-based Strela 1 crane, operated via hand crank by EV1, who will then join EV2. Main objective is the inspection of the Soyuz TMA-12 spacecraft at its first separation plane (Plane I) followed by removal of one pyrobolt for return to Earth. Before removing the separation bolt, EV2 will install temporary protective covers on the spacecraft’s DPO attitude control thrusters, open up the MLI (Multi-Layer Insulation), demate the pyrobolt’s electrical connector and cut the wiretie between the pyrobolts. MLI will then be reattached and the covers removed again. If enough time remains after the Soyuz activity, the spacewalkers will also install a docking target (for the new MLM/Multipurpose Laboratory Module) in the SM PkhO (Transfer Compartment) area for zenith port docking. [Background: Before their separation, Descent Module (SA) & Instrumentation Compartment (AO) are connected by five locks, each “zamok” containing two pyrobolts with individual electrical connection, of which only one needs to fire to release the lock. The locks are equally spaced around the 360-degree circumference of the separation plane, i.e., 72 deg apart. After the five locks have released, five pushers, also equally spaced, separate the two modules by pre-loaded spring force. Each lock and pusher is located at the apex (node) of two triangularly arranged aluminum pipes, i.e., altogether 20 inclined pipes, which make up the open truss structure between the modules.]

EVA-20 Timeline Preview (preliminary): Orlan EVA-20 by Volkov & Kononenko is scheduled to begin on 7/15 at ~1:14m EDT (DC1 EV hatch open), to last an estimated 5 hrs 30 min. Part of the EVA will be supported by the DC1-based Strela 1 crane, operated via hand crank. Main objectives of EVA-20 are –

  • Closeout ops on Soyuz TMA-12 Plane I inspection/pyrobolt removal (if not completed during EVA-20A);
  • Installation of a new docking target on SM PkhO (Transfer Compartment) for zenith port docking of the MLM (if not completed during EVA-20A);
  • Post-installation photography of the new docking target;
  • Inspection of mounting holes for an adapter of a Kurs antenna (4AO-VKA) on PkhO-RO (SM Working Compartment, small diameter section) for MLM;
  • Transferring one “Yakor” foot restraint (of two) from the DC1 EVA ladder to the SM and installing it in an attachment socket at a PkhO handrail (two Yakors were installed on DC1 ladder during EVA-17A on 2/22/07);
  • Installing the VSPLESK (“Burst”) science payload on a handrail at SM RO (large diameter section); and
  • Removing the BIORISK-MSN (BIO-2) experiment container 1 (of three) from the DC1 for return to the station (BIORISK-MSN, with three containers, was installed during EVA-19 on 6/6/07).

Uncrewed Station Ops Planning: One of the contingencies associated with the Orlan EVA-20A next week (and EVA-20 on 7/15) could require uncrewed station operation for some time. As per IMMT (ISS Mission Management Team) decision this morning, SM-to-ATV hatches are to remain open during this period (to be reassessed for EVA-20). ATV batteries are non-rechargeable, thus, time for ATV to be on independent power is very limited. Most of the USOS preparations for the uncrewed period will be done by Gregory Chamitoff next Wednesday (7/9), such as transferring selected hardware to the RS (Russian Segment), setting up a PCS (Portable Computer System) laptop in the FGB as backup to the PCS in the RS, powering down ham/amateur radio equipment, reconfiguring some LAN software (NetMeeting, KFX), powering down the COL PWS (Columbus Orbital Laboratory Portable Workstation) laptop, and closing selected hatches. On 7/10, the IATCS (Internal Thermal Control System) will be configured as usual for uncrewed ops, some racks will be jumpered to the LTL (Low Temperature Loop) in case they need cooling, and some remaining hatches will be closed, before Gregory moves to the Soyuz Descent Module (SA) and closes the hatch between it and the Orbital Module (BO).

CEO (Crew Earth Observations) photo targets uplinked for today were Mt. Kilimanjaro, Kenya (this famous peak, at 19,340 feet, is Africa’s highest mountain and is located in northeastern Tanzania near the border with Kenya. CEO specialists are monitoring the tiny ice fields near the summit that have noticeably diminished during the twentieth century. On this pass, just after midday local time, there may have been low clouds in the surrounding plains as ISS approached the area from the NW. Looking well right of track for the high volcanic crater with patches of ice and snow and using the long lens settings for detail), S. Mozambique (the southern portion of this African nation is undergoing rapid land use change as mineral exploration is driving the construction of new infrastructure. This ISS fair-weather pass in mid-afternoon offered the opportunity to acquire good baseline, contextual views of the southwestern part the target area. There are very few strong visual cues for this area, so as ISS approached the target area from the NW and left the mountain regions of South Africa, Greg was to try for near nadir views in the plains until reaching the coast, then map the coastal area), and Georgia Coastal Ecosystems (the coastal wetlands and barrier islands of Georgia are designated a Long Term Ecological Research [LTER] site. ISS had a fine nadir pass over this target in fair weather conditions, just before noon local time. As the station approached from the NW, the FE-2 was to use the long lens setting to map, in detail, the coastal features northward from the urban areas of Jacksonville, Florida).

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

ISS Orbit (as of this morning, 9:01am EDT [= epoch]):
Mean altitude — 344.8 km
Apogee height — 350.5 km
Perigee height — 339.1 km
Period — 91.43 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0008498
Solar Beta Angle — 5.9 deg (magnitude increasing)
Orbits per 24-hr. day — 15.75
Mean altitude loss in the last 24 hours — 48 m
Revolutions since FGB/Zarya launch (Nov. 98) — 55110

Significant Events Ahead (all dates Eastern Time, some changes possible.):
07/10/08 — Russian EVA-20A (2:18pm)
07/15/08 — Russian EVA-20 (1:14pm)
07/23/08 — ATV1 reboost (tent.)
09/05/08 — ATV1 undocking, from SM aft port (loiter until ~9/25 for nighttime reentry/observation)
09/09/08 — Progress M-64/29P undocking, from FGB nadir (may move to 8/30)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft port)
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
1QTR CY09 — STS-119/Discovery/15A launch – S6 truss segment
03/25/09 — Soyuz TMA-14/18S launch
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking, May ’09)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).

SpaceRef staff editor.